# TreeCat: Standalone Catalog Engine for Large Data Systems

### v1

@@ 4.2. Basic API and Query Language
Query Language: TreeCat’s query language specifies a hierarchical path navigation. It is similar to XPath (Clark et al., 1999), but is much more constrained as it only allows top-down navigation of each context object’s children whereas XPath allows navigation along one of multiple different axes (children, descendants, following, parent etc.) at each step. A query takes form of a path expression, a sequence of $n$ predicates, each of which is evaluated against the children of context objects that satisfy the preceding path subexpression. A predicate is either a wildcard character or an expression (enclosed with [][][ ]) that may be evaluated against the object id or object properties. We extend the example in Figure 2 to construct an example query. A query that asks for files and database blocks containing sales data for clothes in Asia can be constructed as follows:
/[obj_id = "retail"]/[name = "Sales"]/[region = "Asia"]/ [category = "clothes"]/*
The first predicate, [obj_id = "retail"], is evaluated against all objects under the root object, /. Then, the next predicate, [name = "Sales"], is evaluated against the children of the retail database object, which has an object id equal to "retail". The rest of the path expression is evaluated in the same manner as [region = "Asia"] is evaluated against the children of the Sales table object, etc.
@@
    Add: Add operation adds a new object of the given path and initializes it with the given value. The operation has a precondition that the parent object exists, but an object with the same path does not.
@@
    Update: Update operation updates the object of the given path with the given value. If the object does not exist, a new object with the provided value is created.
@@
    Remove: Remove operation removes the object of the given path and all of its descendants.
@@
We describe the merge operation in more detail and highlight its importance. Merge operation is useful when objects are frequently updated by read-modify-write operations (e.g., counters), which may result in many read-write conflicts. An important use case in TreeCat is applying changes in statistics that propagate up the hierarchy as a result of any changes in the low level objects, which can result in contention on high level statistics objects. We present an example of an update to table statistics that follows data insert:
Original Value: {...size:1487...min:3...}
Delta: {...size:{op:+,val:124}...min:{op:min,val:0}...}
Final Value: {...size:1611...min:0...}
@@
Catalog API: We show how catalog API can be built on top of TreeCat with an example from the Spark catalog API we implemented for the experimental evaluation. Each operation takes in an optional TreeTxn, which keeps track of the current transaction session. The following is a simplified snippet of $getTable()$ function, which returns the table information, given a database name and table name.
def getTable(db: String, table: String,
  txn: Option[TreeTxn] = None): CatalogTable = {
  val requestBuilder = queryRequestBuilder(
    f"/[obj_id=\"\$db\" and obj_type=\"database\"]"
    + f"/[obj_id=\"\$table\" and obj_type=\"table\"]")
  setTxnId(requestBuilder, txn)
  val queryResponses = catalogStub
    .executeQuery(requestBuilder.build())
The function constructs a query request with appropriate vid or txn_id if txn is defined. Then, it sends it to the catalog server by a $executeQuery()$ RPC (Remote Procedure Call). queryResponses is a stream of buffers filled with BSON documents, which are deserialized to in-memory CatalogTable object.
@@ 5.1. High Level Architecture
TreeCat is designed as a single-server database engine that handles incoming requests from remote client systems by the gRPC interface. It is written in approximately 12000 lines of C++ code, excluding any third-party libraries. Key components include:
@@
Figure 4. High level architecture of TreeCat
@@ 5.2. Backend Server
We use gRPC (gRPC Authors, 2024), a Remote Procedure Call (RPC) framework developed by Google, to provide the APIs listed in the previous section. gRPC is language-agnostic and offers portability across different languages. While REST API is another option with good portability, gRPC provides better performance due to its use of pre-compiled serialization and deserialization routines and support for server-side streaming, especially for large data loads (it is not uncommon for reads in this context to return MBs of data).
@@ 5.3. Storage Engine
Numerous options exist for full-fledged database systems, ranging from RDBMS with XML (Shanmugasundaram et al., 2001; Tatarinov et al., 2002; O’Neil et al., 2004) and graph (ten Wolde et al., 2023) extensions, to graph databases (Neo4j, 2024; NebbulaGraph, 2024; HyperGraphDB, 2024; ArangoDB, 2024; JanusGraph, 2024; Dubey et al., 2016; Zhu et al., [n.d.]; De Leo and Boncz, 2021; Fuchs et al., 2022), and document stores (MongoDB, 2024; CouchDB, 2024; ArangoDB, 2024). However, meeting all the requirements, especially (2) explicit versioning and (3) fine-grained concurrency control, proves to be difficult, as these systems are designed and implemented for more general use cases. The concurrency control mechanism is especially problematic as general schemes adopted by these systems are not specifically designed to handle hierarchical data. The graph or document databases also don’t typically support the fine-grained concurrency control needed to support updates to portions of the hierarchical metadata.
@@
After careful consideration, we built TreeCat on top of RocksDB, a write-optimized key-value store that is widely used as the storage engine of many systems. Inspired by file systems built on top of write-optimized key-value stores (Jannen et al., 2015; Shetty et al., 2013; Esmet et al., 2012), we adopt a storage layout where objects are sorted by the full object path, first by the path depth, then lexicographically. This layout allows fast listing of sibling objects by performing a range scan on the common prefix. It has advantages over storing the parent-child relation as a separate set of edges, which incurs the overhead of joining the objects (vertices) with edges during traversals. Although each range scan incurs a seek cost, the expense is amortized since objects typically have multiple children; as we show later, query cost is bounded by the query’s selectivity because most irrelevant objects are pruned early. This layout is also more suited for versioning and transactional updates compared to the scheme where an object’s children are stored together as a single list. With the latter, installing a new object version forces installation of an updated version of the parent’s children list, potentially cascading changes to the root and causing read–write conflicts that severely limit concurrency under serializable isolation. This also imposes a fundamental limitation of Iceberg for supporting multi-table transactions.
@@
Object values are stored in BSON format (Commons, [n.d.]), a binary JSON format that originated from MongoDB, is queriable, and can easily be converted to JSON format. While other efficient JSON-compatible binary formats exist (e.g., PostgreSQL’s JSONB), we chose BSON for its robust library support and rich feature set, which enables faster system development. We use two separate storage systems, InnerObjectStore and LeafObjectStore, which store non-leaf objects and leaf objects respectively. LeafObjectStore stores one object value for each path and the $vid$ range during which the object is visible. Unlike leaf objects, multiple past versions of non-leaf objects are persisted. For non-leaf objects, we adopt the “time-travel” storage scheme (Wu et al., 2017) where the most recent version is stored in the main SnapshotStore and the chain of past versions are stored in a separate DeltaStore, as illustrated in Figure 5. SnapshotStore has the following mapping:
@@
where $cur\_vid$ is the $vid$ from which the current snapshot is valid, and $delta\_vid$ is the $start\_vid$ of the previous object value. DeltaStore has the following mapping:
@@
where the object value is visible between $start\_vid$ and $end\_vid$. In DeltaStore, the objects are sorted by the object path, then by $start\_vid$ in descending order, so past object values can be traversed through a range scan. Retrieving the object value of a non-leaf object involves first checking the SnapshotStore for the current value and traversing the past values in DeltaStore, if necessary. Retrieving past versions incurs the cost of version traversal, but we expect most operations will access the most recent version. LeafObjectStore has the following mapping:
@@
where the object is visible between $create\_vid$ and $tombstone\_vid$. Because leaf objects can have multiple parents, there could be a layer of indirection where the path could map to the primary path, which, in turn, maps to the object value.
@@ 5.4. Query Execution
A read query is executed as a sequence of physical operators analogous to the correlated join operator from RDBMSs (also called CROSS APPLY or LATERAL JOIN), which passes each value from the outer query to a user-defined function or a subquery and evaluates it. Unlike other standard join operators, the correlated join operator allows for early pruning of objects in the inner query by only scanning objects that join with objects that satisfy the outer query. We use this mechanism to avoid scanning large number of leaf objects whose ancestors do not satisfy the preceding predicates in the path expression. For example, while executing /[obj_id = "retail"]/[name = "Sales"]/*/*/*, file objects that belong to the Customer table need not be scanned.
Given a path expression of depth $n$, the execution plan is constructed as a single chain of $n$ execution nodes, each with the corresponding predicate, a fixed-sized output buffer, a pointer to its child node, and handles to the object store. The query plan follows a standard batch iterator model where each node invokes $next()$ on its child node, retrieving a batch of context object paths. For each context object path, the executor scans and evaluates the predicate against every child object visible in the $read\_vid$. If the object satisfies the predicate, its path (or its value if the last execution node) is appended to the output buffer. The execution of a simple example query is illustrated in Figure 6. The planner currently applies one optimization rule where the lower and/or upper bounds of the object store iterator are set if the predicate is only satisfied by a range of object ids (e.g. [1 < obj_id < 100]).
@@
$n_{scan}$ $=f+s_{1}f^{2}+...+\prod_{i=1}^{d-1}s_{i}f^{d}$
$\leq\frac{s^{d}f^{d+1}}{sf-1}\text{ where }s=\max_{i=1}^{d-1}s_{i}$
$\leq s^{d}f^{d+1}\text{ when }sf\geq 2$
The number of seek operations is equal to the number of scan operations of the preceding path sub-expression and can be computed to be $s^{d-1}f^{d}$ in a similar manner. Because there is no closed form formula for $f$ in terms of $|T|$, we approximate $f$ to $|T|^{1/h}$. Finally, the total cost can be approximated to $s^{d-1}\cdot|T|^{d/h}(s\cdot|T|^{1/h}\cdot c_{scan}+c_{seek})$ where $c_{scan}$ and $c_{seek}$ represent the cost of a single scan operation and that of a seek operation, respectively. We observe that the multiplier effect of selectivity $s$ effectively bounds the number of scan and seek operations necessary for executing the given query, which aligns with the intuition about “early pruning” of objects. This contrasts with more popular join operators in the relational model, such as sort-merge join or hash join that require at least one scan of the entire table, which can be costly for a large table size (analogously tree size).
@@ 5.5. Version Control Operations
The system supports globally consistent chronological versioning of the entire catalog data, using monotone increasing $vid$ associated with every transaction. $snapshot()$ operation is implemented by maintaining an additional mapping between the snapshot name and the $vid$, which can be used to retrieve the correct version. $clone()$ operation copies the source object and all of its descendants (except for the leaf objects) at some $vid$ to the destination path. For cloned leaf objects, the system adds a mapping from the new destination paths to the primary paths, so they are shared between the source and destination parents, saving the storage costs of duplication.
@@ 6.1. Discussion
We first discuss the background and primary motivations, followed by an in-depth description of the implementation details. We adopt the MVOCC scheme where read-only transactions operate on a versioned snapshot of the data, and read-write transactions follow a protocol similar to the standard OCC mechanism. Because we implemented MVOCC on top of the existing version system, $vid$ is synonymous with $timestamp$ as commonly discussed in the concurrency control literature. For read-only transactions, we follow the common practice of assigning a $read\_vid$, the version up to which data are safe for read, at the start of the transaction, and executing queries against the versions of objects visible at the assigned $read\_vid$. Since this aspect of the mechanism aligns with other existing systems, our main focus is on the OCC, which utilizes a predicate-based method, often overlooked by standard schemes.
@@
The important feature of TreeCat is that read queries involve predicate read operations. Unlike simple key-value read and write operations, predicate read operations can introduce predicate dependencies (Clark et al., 2024), which are distinguished from the item dependencies (write-read, write-write, and read-write) that are commonly discussed in the concurrency control literature. The infamous phantom read anomaly is one example of a predicate anomaly that can occur if predicate dependencies are not handled correctly by the concurrency control mechanism (Eswaran et al., 1976). Unlike item dependencies—which are based on conflicting operations on the same item—predicate dependencies occur when an update alters whether an object satisfies a given predicate. Specifically, if a write operation’s before-image satisfies the predicate of a predicate read operation but its after-image does not, or vice versa, a predicate dependency exists between the two operations (Adya et al., 2000). Item-based approaches, including record locks in 2PL (2 phase locking), read and write sets in OCC, and time stamps on individual items in TO (timestamp ordering), lack semantics for predicate reads, and cannot handle predicate dependencies without additional mechanisms.
@@
There are predicate-based methods that can precisely detect any predicate dependency. One scheme is to re-execute the scan operations to observe any changes in the visible objects (Larson et al., 2011). While precise, re-executing scans can be costly for large ranges. Precision locking (Jordan et al., 1981) is a form of predicate locking that evaluates the predicate directly against posted write operations. This technique detects predicate dependencies accurately and avoids the satisfiability test problem of conventional predicate locking. But in TreeCat, a read query is correlated whence an object satisfies the entire path expression only if all of its ancestors also satisfy the corresponding path sub-expressions. Hence, a naive implementation that only considers the predicates in a given query can, in fact, result in false conflicts.
@@
In our scheme, we combine scan range locking with precision locking, leveraging the strengths of both methods. We first apply optimistic prefix-based scan range locking to detect candidate conflicts, i.e., write operations that fall within the scan range, but may not necessarily cause conflicts. Then, we use optimistic precision locking to confirm whether the candidate is actually a conflict. This scheme achieves the granularity of precision locking while improving efficiency by pruning unnecessary predicate evaluations with scan range locking. In addition, we make use of an in-memory index structure to quickly locate only the log records that fall in the scan range, rather than applying precision locking on the entire set of log records.

### v2

@@ 3.2. Query Language
TreeCat’s query language is based on top-down path navigation, where a query takes the form of a path expression, a sequence of $n$ predicates, each of which is evaluated against the children of context objects that satisfy the preceding path subexpression. A predicate is either a wildcard character or an expression (enclosed with $[]$) that may be evaluated against the object id or object properties. We extend the example in Figure 3 to demonstrate how client systems can utilize TreeCat. Consider how a relational database system would execute the following simple query.
SELECT count(*)
FROM retail.sales
WHERE region = ’US’ AND date > ’2025-01-01’
AND price > 5;
After parsing the query, the system has to semantically validate it. First, it retrieves the metadata of the Sales table from TreeCat, using the following metadata query.
/[obj_id=’retail’]/[obj_id=’sales’
    and obj_type=’table’]
A nonempty query result also validates that the table with the specified database name and table name indeed exists. After the table metadata is retrieved, the system uses it to further check whether column names and types are valid, etc. During the query planning phase, the paths and statistics of the data files that must be scanned during the query execution have to be retrieved. The following query is submitted to retrieve the file-level metadata that satisfy the appropriate predicate filters.
/[obj_id=’retail’]/[obj_id=’sales’ and
    obj_type=’table’]/[part_val=’US’]
    /[part_val > ’2025-01-01’]/[stats.price.min > 5]
The physical plan can finally be executed, once it is constructed.
We briefly discuss related work, including the APIs of existing standalone catalogs and query languages of similar, more general-purpose database systems. Standalone catalogs, such as HMS and lakehouse storage formats, provide basic RPC routines, such as getTable() and getPartitions(), via Thrift or the REST API for retrieving high-level catalog objects. Retrieving file-level metadata of a given table involves iterating the list of file paths via an iterator interface with an optional predicate filter, rather than a unified query language like TreeCat. Meanwhile, the query languages of general graph databases, including Cypher (Neo4j, 2025a), Gremlin (Foundation, 2025d), and GSQL (TigerGraph, 2025), and those of XML databases, including XPath (Clark et al., 1999) and XQuery (W3C, 2017), are more powerful than TreeCat’s query language and can express a wider range of path queries. However, advanced semantics makes efficient implementation of a strong transaction isolation level very difficult. To our knowledge, there are currently no efficient graph database or XML database that fully support serializable isolation level. In our design, we sacrifice the query expressivity, so we can guarantee serializability using techniques that are discussed in Section 5.
3.3. System API
Basic API: Next, we list a basic API that TreeCat supports.
  • startTransaction(mode): Starts a new transaction in the specified mode. If the mode is read-only, it returns the latest vidvid at which a snapshot of the catalog can be read. If the mode is read-write, it creates a new transaction session and returns transaction_idtransaction\_id and read_vidread\_vid.
  • executeQuery(query, optional transaction_id, optional vid): Executes a read query on the catalog using the query language defined above.
  • commit(transaction_id, write_set): Validates and commits the transaction identified by transaction_idtransaction\_id, applying modifications specified by the given write_setwrite\_set.
  • snapshot(snapshot_name, vid): Creates an immutable snapshot of the catalog data at vidvid and names it snapshot_namesnapshot\_name.
  • clone(src_path, dest_path, optional vid): Clones/copies the object at src_pathsrc\_path and all of its descendants to the dest_pathdest\_path. Note that this is a metadata-level operation, and does not involve cloning the physical objects.
@@ 3.3. System API
    Add: a new object of the given path and initialize it with the given value. The operation has a precondition that the parent object exists, but an object with the same path does not.
@@
    Update: the object with the given path with the given value. A new object is created if the path does not exist.
@@
    Remove: the object with the given path and all its descendants.
@@
We describe the merge operation in more detail and highlight its importance. Merge operation is useful when objects are frequently updated by commutative read-modify-write operations (e.g., counters), which may result in many conflicts. An important use case in TreeCat is applying changes in statistics that propagate up the hierarchy, which can result in contention on high-level statistics objects. We present an example of an update to table statistics that follows data insert:
Original Value: {...size:1487..min:3...}
Delta: {...size:{op:+,val:124}..min:{op:min,val:0}...}
Final Value: {...size:1611..min:0...}
@@ 4.1. High Level Architecture
TreeCat is designed as a single-server database engine that handles incoming requests from remote client systems via the gRPC interface. Key components include:
@@
This architecture allows a single TreeCat instance (possibly replicated, see Section 7) to handle the metadata for several different logical databases, and concurrently serve many execution engines while ensuring metadata consistency. Multiple TreeCat instances could be spun to serve different logical databases for administrative autonomy; however, any logical database (i.e., a collection of logically related tables or datasets) should ideally be served by a single TreeCat instance so that cross-table consistency can be enforced.
Figure 4. High-level architecture of TreeCat
@@ 4.2. Backend Server
We use gRPC (gRPC Authors, 2025), a Remote Procedure Call (RPC) framework developed by Google, to provide the APIs listed in the previous section. gRPC is language-agnostic and offers portability across different languages. While REST API is another option with good portability, gRPC provides better performance due to its use of pre-compiled serialization and deserialization routines and support for server-side streaming, especially for large data loads (it is not uncommon for reads in this context to return MBs of data).
@@ 4.3. Storage Engine
Numerous options exist for full-fledged database systems, ranging from RDBMS with XML (Shanmugasundaram et al., 2001; Tatarinov et al., 2002; O’Neil et al., 2004) and graph (ten Wolde et al., 2023) extensions, to graph databases (Neo4j, 2025b; NebbulaGraph, 2025; HyperGraphDB, 2025; ArangoDB, 2025; JanusGraph, 2025; Dubey et al., 2016; Zhu et al., 2020; De Leo and Boncz, 2021; Fuchs et al., 2022), and document stores (MongoDB, 2025; CouchDB, 2025; ArangoDB, 2025). However, meeting all the requirements, especially (2) explicit versioning and (3) fine-grained concurrency control, proves to be difficult, as these systems are designed and implemented for more general use cases. As discussed before, lack of efficient support for serializable isolation level is especially problematic for graph databases and document stores.
@@
After careful consideration, we built TreeCat on top of RocksDB, a write-optimized key-value store. Inspired by file systems built on top of write-optimized key-value stores (Jannen et al., 2015; Shetty et al., 2013; Esmet et al., 2012), we adopt a storage layout where objects are sorted by the full object path, first by the path depth, then lexicographically. This layout allows fast listing of sibling objects by performing a range scan on the common prefix. It has advantages over storing the parent-child relation as a separate set of edges, which incurs the overhead of joining the objects (vertices) with edges during traversals. Although each range scan incurs a seek cost, the expense is amortized since objects typically have multiple children; as we show later, query cost is bounded by the query’s selectivity because most irrelevant objects are pruned early. This layout is also more suited for versioning and transactional updates compared to the scheme where an object’s children are stored together as a single list. With the latter, installing a new object version forces installation of an updated version of the parent’s children list, potentially cascading changes to the root and causing read–write conflicts that severely limit concurrency under serializable isolation. This also imposes a fundamental limitation of Iceberg for supporting multi-table transactions.
@@
Object values are stored in BSON format (Commons, 2025), a binary JSON format that originated from MongoDB, is queriable, and can easily be converted to JSON format. While other efficient JSON-compatible binary formats exist (e.g., PostgreSQL’s JSONB), we chose BSON for its robust library support and rich feature set, which enables faster system development. We use two separate storage systems, InnerObjectStore and LeafObjectStore, which store non-leaf objects and leaf objects respectively. LeafObjectStore stores one object value for each path and the vidvid range during which the object is visible. Unlike leaf objects, multiple past versions of non-leaf objects are persisted. For non-leaf objects, we adopt the “time-travel” storage scheme (Wu et al., 2017) where the most recent version is stored in the main SnapshotStore and the chain of past versions are stored in a separate DeltaStore, as illustrated in Figure 5. SnapshotStore has the following mapping:
@@
where cur_vidcur\_vid is the vidvid from which the current snapshot is valid, and delta_viddelta\_vid is the start_vidstart\_vid of the previous object value. DeltaStore has the following mapping:
@@
where the object value is visible between start_vidstart\_vid and end_vidend\_vid. In DeltaStore, the objects are sorted by the object path, then by start_vidstart\_vid in descending order, so past object values can be traversed through a range scan. Retrieving the object value of a non-leaf object involves first checking the SnapshotStore for the current value and traversing the past values in DeltaStore, if necessary. Retrieving past versions incurs the cost of version traversal, but we expect most operations will access the most recent version. LeafObjectStore has the following mapping:
@@
where the object is visible between create_vidcreate\_vid and tombstone_vidtombstone\_vid. Because leaf objects can have multiple parents, there could be a layer of indirection where the path could map to the primary path, which, in turn, maps to the object value.
@@ 4.4. Query Execution
A read query is executed as a sequence of physical operators analogous to the correlated join operator from RDBMSs (also called CROSS APPLY or LATERAL JOIN), which passes each value from the outer query to a user-defined function or a subquery and evaluates it. Unlike other standard join operators, the correlated join operator allows for early pruning of objects in the inner query by only scanning objects that join with objects that satisfy the outer query. We use this mechanism to avoid scanning large number of leaf objects whose ancestors do not satisfy the preceding predicates in the path expression. For example, while executing /[obj_id = ‘retail’]/[obj_id = ‘sales’]/*/*/*, file objects that belong to the Customer table need not be scanned.
Given a path expression of depth $n$, the execution plan is constructed as a single chain of $n$ execution nodes, each with the corresponding predicate, a fixed-sized output buffer, a pointer to its child node, and handles to the object store. The query plan follows a standard batch iterator model where each node invokes $next()$ on its child node, retrieving a batch of context object paths. For each context object path, the executor scans and evaluates the predicate against every child object visible in the read_vidread\_vid. If the object satisfies the predicate, its path (or its value if the last execution node) is appended to the output buffer. The execution of a simple example query is illustrated in Figure 6. The planner currently applies one optimization rule where the lower and/or upper bounds of the object store iterator are set if the predicate is only satisfied by a range of object ids (e.g. [1 < obj_id < 100]).
@@
$n_{scan}$ $=f+s_{1}f^{2}+...+\prod_{i=1}^{d-1}s_{i}f^{d}\leq\frac{s^{d}f^{d+1}}{sf-1}\leq s^{d}f^{d+1}$
where $s=\max_{i=1}^{d-1}s_{i}$ and $sf\geq 2$. The number of seek operations is equal to the number of scan operations of the preceding path sub-expression and can be computed to be $s^{d-1}f^{d}$ in a similar manner. Because there is no closed form formula for $f$ in terms of $|T|$, we approximate $f$ to $|T|^{1/h}$. Finally, the total cost can be approximated to $s^{d-1}\cdot|T|^{d/h}(s\cdot|T|^{1/h}\cdot c_{scan}+c_{seek})$ where $c_{scan}$ and $c_{seek}$ represent the cost of a single scan operation and that of a seek operation, respectively. We observe that the multiplier effect of selectivity $s$ effectively bounds the number of scan and seek operations necessary for executing the given query, which aligns with the intuition about “early pruning” of objects. This contrasts with more popular join operators, such as sort-merge join or hash join that require at least one scan of the entire table, which can be costly for a large table size (analogously tree size).
@@ 4.5. Version Control Operations
The system supports globally consistent chronological versioning of the entire catalog data, using monotone increasing vidvid associated with every transaction. $snapshot()$ operation is implemented by maintaining an additional mapping between the snapshot name and the vidvid, which can be used to retrieve the correct version. $clone()$ operation copies the source object and all of its descendants (except for the leaf objects) at some vidvid to the destination path. For cloned leaf objects, the system adds a mapping from the new destination paths to the primary paths, so they are shared between the source and destination parents, saving the storage costs of duplication.
@@ 5.1. Discussion
We first discuss the background and primary motivations, followed by an in-depth description of the implementation details. We adopt the MVOCC scheme where read-only transactions operate on a versioned snapshot of the data, and read-write transactions follow a protocol similar to the standard OCC mechanism. Because we implemented MVOCC on top of the existing version system, vidvid is synonymous with timestamptimestamp as commonly discussed in the concurrency control literature. For read-only transactions, we follow the common practice of assigning a read_vidread\_vid, the version up to which data are safe for read, at the start of the transaction, and executing queries against the versions of objects visible at the assigned read_vidread\_vid. Since this aspect of the mechanism aligns with other existing systems, our main focus is on the OCC, which utilizes a predicate-based method, often overlooked by standard schemes.
@@
The important feature of TreeCat is that read queries involve predicate read operations. Unlike simple key-value read and write operations, predicate read operations can introduce predicate dependencies (Adya et al., 2000; Clark et al., 2024), which are distinguished from the item dependencies (write-read, write-write, and read-write) that are commonly discussed in the concurrency control literature. The infamous phantom read anomaly is one example of a predicate anomaly that can occur if predicate dependencies are not handled correctly by the concurrency control mechanism (Eswaran et al., 1976). Unlike item dependencies—which are based on conflicting operations on the same item—predicate dependencies occur when an update alters whether an object satisfies a given predicate. Specifically, if a write operation’s before-image satisfies the predicate of a predicate read operation but its after-image does not, or vice versa, a predicate dependency exists between the two operations (Adya et al., 2000). Item-based approaches, including record locks in 2PL (2-Phase Locking), read and write sets in OCC, and time stamps on individual items in TO (timestamp ordering), lack semantics for predicate reads, and cannot handle predicate dependencies without additional mechanisms.
@@
There are predicate-based methods that can precisely detect any predicate dependency. One scheme is to re-execute the scan operations to observe any changes in the visible objects (Larson et al., 2011). While precise, re-executing scans can be costly for large ranges. Precision locking (Jordan et al., 1981) is a form of predicate locking that evaluates the predicate directly against posted write operations. This technique detects predicate dependencies accurately and avoids the satisfiability test problem of conventional predicate locking. But in TreeCat, a read query is correlated whence an object satisfies the entire path expression only if all of its ancestors also satisfy the corresponding path sub-expressions. A naive implementation that only considers the predicates in a given query can, in fact, result in false conflicts.
@@
In our scheme, we combine scan range locking with precision locking, leveraging the strengths of both methods. We first apply optimistic prefix-based scan range locking to detect candidate conflicts, i.e., write operations that fall within the scan range, but may not necessarily cause conflicts. Then, we use optimistic precision locking to confirm whether the candidate is actually a conflict. This scheme achieves the granularity of precision locking while improving efficiency by pruning unnecessary predicate evaluations with scan range locking. In addition, we make use of an in-memory index structure to quickly locate only the log records that fall in the scan range, rather than applying precision locking on the entire set.
5.2. Implementation
The concurrency control is handled by the TransactionManager, which keeps track of read_vidread\_vid, the latest version up to which, the catalog can be read safely, and commit_vidcommit\_vid, which is the monotone increasing vid assigned to a transaction before final commit. Key in-memory data structures include:
  1. (1)
    ScanSetScanSet: Maintained by each transaction. Stores pairs of the scan range (in the form of the context object path) and the predicate of the predicate read operations executed as part of the transaction.
  2. (2)
    LogBufferLogBuffer: Maintained by each transaction. Stores before- and after-images of all recent write operations.
  3. (3)
    VersionMapVersionMap: Used by TransactionManager for validation. Maps object path to the most recent vidvid of its children.
  4. (4)
    LogIndexLogIndex: Used by TransactionManager for validation. Index on the $LogBuffer$, mapping pairs of object path and vidvid to the log records. The index is implemented using a lock-free skip list and is sorted by the parent path, then vidvid.
When $startTransaction()$ is called, a new transaction session is created and read_vidread\_vid is loaded as the vid at which all read queries are executed in the session. During query execution, pairs of predicates and parent paths (prefix) of objects scanned for predicate evaluation are added to the $ScanSet$. Also, a private copy of the write set is constructed on the client side. When the transaction is submitted for the commit process, the write set is sent to the server and undergoes preprocessing, where it is copied to the $LogBuffer$ and preconditions are checked. If the transaction passes these checks, the transaction is scheduled for the main commit process.
The main commit process consists of two phases: validation and batch write. For each phase, there is a dedicated thread pool of a fixed size, with each thread handling a single hash partition of the $ScanSet$ and the write set. At the start of the validation, TransactionManager executes a fetch_and_add()fetch\_and\_add() operation on the commit_vidcommit\_vid and assigns it to the transaction. Then, each worker thread validates the scan ranges in the $ScanSet$ against the $VersionMap$ to determine whether any objects scanned by predicate read operations have changed. If potential conflicts are detected, an iterator on $LogIndex$ is set to scan and apply precision locks on corresponding recent writes with vidvid greater than the transaction’s read_vidread\_vid and fall in the scan range. If any conflict is detected through precision locking, the transaction is aborted. After all threads pass validation, they update the $VersionMap$ and $LogIndex$ based on their respective write set partitions for the transaction. The validation phase (omitting the initial vidvid assignment and latter updates to in-memory data structures) proceeds as per Algorithm 1. Note that for merge operation, before- and after-image may have to be constructed as they are not available until the final commit. However, once constructed, the images are cached so they are accessible during the validation of subsequent transactions.
Once validation is complete, the transaction proceeds to the write phase. During this phase, each thread processes its assigned write partition, evaluating the final values of merge operations (commit-time updates) and applying the writes as an atomic batch write operation to the underlying RocksDB instance. After every thread finishes its batch write, WAL records of the transaction are flushed to disk along with those of other transactions via $SyncWAL()$ operation, which calls $fsync()$ on the WAL files. The read_vidread\_vid is also updated, making the committed changes visible for subsequent reads. Transactions are processed one at a time in each phase, but parallelism is achieved via hash-partitioning the workload and pipelining different phases of the commit process.
Old $LogBuffer$s and entries in $VersionMap$ and $LogIndex$ that are no longer needed are asynchronously garbage collected by a background thread. For this purpose, the TransactionManager keeps track of watermarkwatermark, the oldest read_vidread\_vid among the current transactions. After each transaction commit, the watermarkwatermark is updated and all entries in the aforementioned data structures with vidvid below the watermarkwatermark are garbage collected.